A flywheel of mass 15kg and a radius of 0.2m is mounted on a horizontal axle of radius 0.01m and negligible mass A force of 40N is applied tangentially to the axle.Find moment of inertia of the wheel,torgue, angular acceleration and angular velocity at the 10 s




1
Expert's answer
2021-12-14T09:28:14-0500

"I=\frac{m(R-r)^2}2=0.27~kg\cdot m^2,"

"T=FR=8~N\cdot m,"

"\varepsilon=\frac TI=29.6~\frac{rad}{s^2},"

"\omega=\varepsilon t=296~\frac{rad}s."